# Signed Higher-Order Interactions for Brain Disorder Diagnosis via Multi-Channel Transformers

### v2

@@ 5.1 The model
Furthermore, the higher-order organisational patterns identified by HOI-Brain offer novel insights into the diagnosis of brain diseases. Specifically, the quadruplet interactions among CAU.R, HIP.L, PHG.R, and AMYG.L; the quadruplet interactions among OLF.L, HIP.L, PHG.R, and AMYG.L; and the quadruplet interactions among CAU.R, OLF.L, PHG.R, and AMYG.L, which involve the cortico-striatal-olfactory circuit, may serve as potential biomarkers for the early diagnosis of Alzheimer’s disease with p-values <0.05<0.05< 0.05 during the progression from cognitively normal (CN) to mild cognitive impairment (MCI). It is noteworthy that the brain regions identified by HOI-Brain via the analysis of resting-state fMRI data are deemed to be of significant importance in the diagnosis of Alzheimer’s disease (AD) or Autism Spectrum Disorder (ASD). This finding indicates that these brain regions may have sustained varying degrees of damage. However, higher-order interactions exhibited by these regions demonstrate notable differences across various stages of disease progression. This finding is consistent with the conclusions of previous studies (Wang et al., 2019), suggesting a dissociation between higher-order interactions in the brain and the functions of the individual brain regions themselves. Furthermore, observations were made concerning the transition from healthy individuals to those with Alzheimer’s disease (AD) or autism spectrum disorder (ASD). It was noted that the positive higher-order interactions between relevant brain regions gradually weaken, indicating a progressive impairment of higher-order synergistic functions among brain regions. Conversely, the negative higher-order interactions between these brain regions gradually strengthen, reflecting an increasing degree of internal functional disruption in the brain. This phenomenon is consistent with the findings reported in previous literature (Santoro et al., 2024). However, this phenomenon is exactly the opposite in Parkinson’s disease, which may be due to the fact that patients suffering from Parkinson’s disease, due to their motor impairments, are prone to hallucinations. Hallucinations in Parkinson’s disease may be attributable to an excessive influence of higher-order brain regions on early sensory processing. This is consistent with enhanced functional integration between sensory and higher-order networks (Tan et al., 2023).
5.2 Limitation and future work
@@
We acknowledge several limitations of our study, which also serve as inspiration for future research: 1) Ignored the discordant signs of negative: In this paper, we focus solely on the concordant signs of positive effects, while only exploring the influence of discordant signs of negative within the interpretable section. These discordant signs are known to play a critical role in the overall system dynamics, as they may reflect redundant or conflicting information across multiple brain regions. Future work should investigate alternative methods to more explicitly address discordant patterns in the data. 2) Lack of dynamic exploring for higher-order topological features: To mitigate noise in fMRI data, we analyze average statistics across multiple time frames, which helps extract stable low-order and higher-order patterns. However, our current approach does not account for the temporal dependencies in higher-order topological features—i.e., how prior time points influence subsequent ones. Since our model reveals that these higher-order patterns evolve dynamically over time, further investigation into this temporal evolution could provide novel insights into disease progression. This opens up promising avenues for modeling the dynamic nature of brain network topology. 3) Limited modality of experimental data: Currently, our analysis is based solely on fMRI data. While these data effectively demonstrate the importance of higher-order interactions in brain network analysis, other time-series modalities—such as EEG or MEG—remain unexplored. In future work, we plan to extend our framework to multimodal neuroimaging data and develop methods tailored to capturing higher-order interactions across different temporal scales.
@@ 6 Conclusions
In summary, we propose a new framework, HOI-Brain (Higher-Order Interactions in Brain Network), which can accurately capture signed higher-order interactions in brain networks, extract interpretable signed higher-order topological features, and further exploit the information between lower-order and higher-order features for brain disorder diagnosis. We applied HOI-Brain to datasets for Alzheimer’s disease, Parkinson’s disease, and Autism Spectrum Disorder. With the powerful interpretability, HOI-brain not only performs better on classification than 20 baselines, but also detects salient brain regions associated with classification and discovers important higher-order organizations. Overall, our model shows superiority over alternative graph learning and traditional machine learning classification models. By analyzing the attention maps of our multi-channel brain Transformer, our study identifies salient regions of interest (ROIs) and their key interactions—from a whole-brain perspective—to distinguish brain disorders from healthy controls. Additionally, we uncover higher-order organizational patterns associated with specific disease progression stages. Notably, our framework is generalizable to the analysis of other neuroimaging time-series data. Our proposed framework demonstrates the potential of quadruplet or higher-order region interactions—critical advancements for precision medicine, enhanced understanding of neurological disorders, and broader contributions to neuroimaging research.

### v3

@@ 5.1 The model
Furthermore, the higher-order organisational patterns identified by HOI-Brain offer novel insights into the diagnosis of brain diseases. Specifically, the quadruplet interactions among CAU.R, HIP.L, PHG.R, and AMYG.L; the quadruplet interactions among OLF.L, HIP.L, PHG.R, and AMYG.L; and the quadruplet interactions among CAU.R, OLF.L, PHG.R, and AMYG.L, which involve the cortico-striatal-olfactory circuit, may serve as potential biomarkers for the early diagnosis of AD with p-values <0.05<0.05< 0.05 during the progression from CN to MCI. It is noteworthy that the brain regions identified by HOI-Brain via the analysis of resting-state fMRI data are deemed to be of significant importance in the diagnosis of AD or ASD. This finding indicates that these brain regions may have sustained varying degrees of damage. However, HOIs exhibited by these regions demonstrate notable differences across various stages of disease progression. This finding is consistent with the conclusions of previous studies (Wang et al., 2019), suggesting a dissociation between HOIs in the brain and the functions of the individual brain regions themselves. Furthermore, observations were made concerning the transition from healthy individuals to those with AD or ASD. It was noted that the positive HOIs between relevant brain regions gradually weaken, indicating a progressive impairment of higher-order synergistic functions among brain regions. Conversely, the negative HOIs between these brain regions gradually strengthen, reflecting an increasing degree of internal functional disruption in the brain. This phenomenon is consistent with the findings reported in previous literature (Santoro et al., 2024). However, this phenomenon is exactly the opposite in Parkinson’s disease, which may be due to the fact that patients suffering from Parkinson’s disease, due to their motor impairments, are prone to hallucinations. Hallucinations in Parkinson’s disease may be attributable to an excessive influence of higher-order brain regions on early sensory processing. This is consistent with enhanced functional integration between sensory and higher-order networks (Tan et al., 2023).
5.2 Future work
@@
Through our work, we have achieved good accuracy and interpretability in brain disease diagnosis. There are still some aspects worth exploring in the future. In this paper, we focus on the concordant signs of positive effects. However, exploring the influence of discordant signs of negative effects within the interpretable section, we found that these discordant signs may reflect redundant or conflicting information across multiple regions of the brain. Future work may investigate alternative methods to more explicitly address discordant patterns in the data. In addition, our model can reveals that these higher-order patterns evolve dynamically over time, so further investigation into this temporal evolution could provide novel insights into disease progression. This opens up promising avenues for modeling the dynamic nature of brain network topology. Finally, our model demonstrates strong generalizability and can be adapted to other time-series modalities—such as EEG or MEG—enabling further exploration of the importance of HOIs in the brain.
@@ 6 Conclusions
In summary, we propose a new framework, HOI-Brain (Higher-Order Interactions in Brain Network), which can accurately capture signed HOIs in brain networks, extract interpretable signed higher-order topological features, and further exploit the information between lower-order and higher-order features for brain disorder diagnosis. We applied HOI-Brain to datasets for Alzheimer’s disease, Parkinson’s disease, and Autism Spectrum Disorder. With the powerful interpretability, HOI-brain not only performs better on classification than 20 baselines, but also detects salient brain regions associated with classification and discovers important higher-order organizations. Overall, our model shows superiority over alternative graph learning and traditional machine learning classification models. By analyzing the attention maps of our multi-channel brain Transformer, our study identifies salient ROIs and their key interactions—from a whole-brain perspective—to distinguish brain disorders from healthy controls. Additionally, we uncover higher-order organizational patterns associated with specific disease progression stages. Notably, our framework is generalizable to the analysis of other neuroimaging time-series data. Our proposed framework demonstrates the potential of quadruplet or higher-order region interactions—critical advancements for precision medicine, enhanced understanding of neurological disorders, and broader contributions to neuroimaging research.